Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dr. James McDaid, TD Minister for Tourism, Sport and Recreation Department of Tourism, Sport and Recreation Kildare Street Dublin 2

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Sports Capital Grant programme 2002  10th December 2001

Minister,

I write representing the interests of the Dublin University Rifle Club. We are an Olympic Target Shooting club in Trinity College Dublin. We have recently discovered that, despite being available to 17,000 current third-level students, 10,000 current Trinity staff and in excess of 80,000 alumni (a very conservative estimate), we are denied access to the Sports Capital Grant programme on the basis that we are not available to the local community.

We have a long and distinguished history in the sport of Olympic target shooting, including winning Ireland’s first international gold medal in the 10 metre Air Rifle event. The club is run completely by volunters. We do not discriminate on the basis of race, sex, age or religion. We currently have active members of all ages from 17 to 40 and have had members in the past who were active past the age of 60. In fact, the one and only restriction on membership in the club is that the members are members of the Dublin University Central Athletics Club and this restriction is outside our control, being mandated by the College. However, entry to third-level education in Ireland has no restrictions and is available, free from tuition fees and with grant aid from the Department of Education, to anyone who wishes it and is prepared to study.

College and University clubs represent the primary area for young people to begin in what are regarded as minority sports. During their time in third-level education, young people may experiment with a multitude of sports, with which they would have had no contact with prior to then, simply because sports equipment is too expensive, or the sport is considered too esoteric for a local club. Trinity currently has 45 sports clubs - the NUI colleges have similar numbers. All of these clubs are open to all students, staff and alumni of their respective colleges. The numbers of people who may join one of these clubs are staggering - we conservatively estimate that there are approximately 107,000 people between the ages of 17 and 60 who are eligible for membership in DURC today.

This surely qualifies as a community - indeed, there are more people currently in Trinity today than in most towns in Ireland. UCD finds itself in the same situation, as do sports clubs in all third-level institutions.

I also serve on the committee of the National Target Shooting Association of Ireland, the governing body for Olympic rifle shooting in Ireland, and in that capacity was present at the Sports Development Conference in November of this year. At that conference, the message from the Irish Sports Council was that the Department was committed to furthering Sport in Ireland at all levels from local community clubs to High Performance athletes.

Minister, I ask you to show that the Department of Sport’s committment goes beyond mere lip service, to meaningful support of one of the primary areas through which young people today can gain access to sporting facilities - the college sports clubs. I ask that we be given equal status to the sports clubs in local communities in light of the fact that we provide facilities to the largest communities of young people in Ireland today. I ask that we be allowed access to the Sports Capital Grant programme.

Misé le meas,

Mark Dennehy, DURC Secretary